# VinePPO: Unlocking RL Potential For LLM Reasoning Through Refined Credit Assignment

### v1

@@ 4 Accurate Credit Assignment with VinePPO
As outlined in Section 3, a step in the PPO gradient update aims to increase the probability of better-than-average actions while decreasing the probability of those that perform worse—a process quantified by the advantage $A(s_{t},a_{t})$. However, the true advantage is generally unknown and must be estimated, typically by substituting estimates from a value network into Equation 3. As we will elaborate in Section 7, value networks are often inaccurate and result in biased value computation. Fortunately, the language environment as an MDP (Section 3) offers a useful property that allows for unbiased estimation of $V(s_{t})$. Since states are simply concatenated tokens, we can prompt the language model $\pi_{\theta}$ to generate continuations from any intermediate state. This flexibility allows us to explore alternative future paths from arbitrary points in a generation Moreover, recent advancements in LLM inference engines (Kwon et al., 2023; Zheng et al., 2024) have dramatically increased the speed of on-the-fly response generation666up to 5K tokens/second on a single Nvidia A100 GPU for a 7B LLM loaded in bfloat16.. This computational efficiency makes it feasible to conduct fast environment simulation, opening up unique opportunities for RL training of LLMs.VinePPO uses this property and estimates advantage via MC sampling. It only modifies the way advantages are estimated, leaving the rest of the standard PPO pipeline intact (Figure 2).
We start by estimating the true value $V(s_{t})$. Instead of relying on a value network, for any intermediate state $s_{t}$, we sample $K$ independent trajectories $\tau^{k}$’s. The average return across these trajectories serves as the value estimate:
$\hat{V}_{\mathrm{MC}}(s_{t})\coloneqq\frac{1}{K}\sum_{k=1}^{K}R(\tau^{k}),\quad\text{where}\ \tau^{1},\dots,\tau^{K}\sim\pi_{\theta}(\cdot\mid s_{t}).$ (5)
This is a MC estimate of $V(s_{t})=\mathbb{E}\left[R(\tau)\mid s_{0}=s_{t}\right].$ Note that these trajectories are not trained on. Once the value $\hat{V}_{\mathrm{MC}}(s_{t})$ is computed, we estimate the advantages of each action using Equation 3:
$\hat{A}_{\mathrm{MC}}(s_{t},a_{t})\coloneqq r(s_{t},a_{t})+\gamma\hat{V}_{\mathrm{MC}}(s_{t+1})-\hat{V}_{\mathrm{MC}}(s_{t}).$ (6)
For any $K\geq 1$, the policy gradient computed using the advantage estimator $\hat{A}_{\mathrm{MC}}$ is an unbiased estimate of the gradient of expected return ${\bm{g}}_{\mathrm{pg}}$. To enhance the efficiency of $\hat{A}_{\mathrm{MC}}$, we group states within a reasoning step and compute a single advantage, which is assigned to all tokens in that step (examples in Appendix B). This trades off granularity for efficiency, allowing finer resolution with more compute, or coarser estimates with limited resources. The parameter $K$ also offers another trade-off between computational cost (i.e. more MC samples per state) and the variance of the estimator. As shown in Section 6.1, even $K=1$ performs well.
Figure 5: Comparison of the training behavior between VinePPO and PPO. VinePPO demonstrates consistently higher accuracy (as measured on the test set of MATH dataset) throughout the training. Refer to Appendix D for more detailed plots.
In essence, VinePPO is a straightforward modification to the PPO pipeline, altering only the advantage computation. This minimal adjustment allows us to leverage PPO’s benefits while enabling a systematic evaluation of the effect of unbiased advantage estimation and improved credit assignment. In the following sections, we compare various aspects such as task performance, computational efficiency, KL divergence, and robustness to shed light on the nature of these approaches.
@@ 5 Experimental Setup
We conduct our experiments using LLMs that show strong performance on mathematical reasoning: DeepSeekMath 7B (Shao et al., 2024) and RhoMath 1.1B (Lin et al., 2024), both of which have been trained on diverse mathematical and natural language corpora. Having different sized models allows evaluating the effect of scaling. We focus on mathematical reasoning datasets MATH (Hendrycks et al., 2021), consisting of competition-level mathematical problems, and GSM8K (Cobbe et al., 2021), containing simpler grade-school level math word problems. Both datasets are well-established and present a range of difficulty levels that allow for comprehensive evaluation. For each dataset, we finetune the base LLMs on its respective training sets to obtain the initial SFT policy ($\pi_{\mathrm{ref}}$). In all experiments, we employ full-parameter finetuning to allow utilization of models’ full capacity (Sun et al., 2023; Biderman et al., 2024).
Evaluation
We evaluate model performance on the test sets of each dataset, using accuracy (Pass@1) as our primary metric, which measures the correctness of the final answers produced by the models. As our baseline, we adopt the standard PPO framework, as commonly implemented for LLM finetuning (Ouyang et al., 2022; Huang et al., 2024). Additionally, we compare them against RL-free methods that doesn’t have explicit credit assignment mechanisms: RestEM (Singh et al., 2024), a form of Iterative Rejection Finetuning (Yuan et al., 2023; Anthony et al., 2017) and DPO+ (Pal et al., 2024), variant of DPO with strong performance on reasoning tasks. All methods are initialized from the same SFT checkpoint to ensure a fair comparison.
@@
To ensure standard PPO (and its value network) has a healthy training and our evaluation reflects its full potential, we first focus our hyperparameter search on PPO parameters (such as KL penalty coefficient, batch size, minibatch size, GAE $\lambda$, number of epochs per iteration) and apply all well-known techniques and best practices (Huang et al., 2024; Ivison et al., 2024) in PPO tuning (Refer to Section C.2 for the full list). Following previous work (Pal et al., 2024; Singh et al., 2024), we set the task reward $\mathcal{R}$ to be a binary function that only checks final answer against the ground truth. VinePPO borrows the exact same hyperparameters from PPO and only modifies the advantage $A(s_{t},a_{t})$ estimation, keeping the rest of the pipeline unchanged. This allows us to isolate the effect of accurate credit assignment. We found that sampling $K=9$ trajectories in $\hat{V}_{\mathrm{MC}}$ performs well; the effect of varying $K$ is fully analyzed in Section 6.1. For the other baseline, we closely follow the original setup while ensuring consistency in training conditions for a fair comparison. We choose the best checkpoint based on a held-out validation set for all experiments. Full implementation details, including all hyperparameters and training procedures, are provided in Section C.6.
@@ 6 Results
We evaluate the effect of accurate credit assignment on four key measures of model finetuning efficiency and success: task performance, KL divergence, temperature tolerance, and computational efficiency. Our experimental setup is designed to control for and isolate the impact of credit assignment on each of these measures.
Figure 6: Task accuracy as a function of KL divergence during training on the MATH dataset. VinePPO achieves higher accuracy, reflecting more efficient credit assignment and focused updates.
@@ 6.1 Task Performance
VinePPO consistently outperforms standard PPO throughout training (Figure 5) and other baselines (Figure 3). More importantly, its performance gap widens in MATH which is a much more challenging reasoning task. Unlike VinePPO and PPO, DPO+ and RestEM lacks any explicit mechanisms for credit assignment, opting instead to finetune the model on the full trajectory. Our experiments show that these RL-free methods lags behind both PPO-based methods. For RestEM, the absence of targeted credit assignments likely leads to overfitting (Section C.5).
To assess the impact of $K$, the number of MC samples used to estimate the value, we run an ablation on RhoMath 1.1B, varying $K$ from 1 to 3 and then to 9. As shown in Figure 4, VinePPO demonstrates improved performance with higher $K$ values, as more MC samples reduce the variance of the $\hat{A}_{\mathrm{MC}}$ estimator. Notably, increasing $K$ provides a reliable approach to leveraging additional computational resources for better performance.
6.2 KL Divergence
Figure 8: Test set accuracy during training with higher temperature presented for DeepSeekMath 7B and MATH dataset. VinePPO can tolerate higher temperatures.
The RL objective (Equation 1) balances maximizing task performance while constraining deviations from the initial policy $\pi_{\mathrm{ref}}$, measured by KL divergence. We analyze how VinePPO and PPO navigate this trade-off by plotting task accuracy against KL divergence $\mathrm{KL}[\pi_{\theta}\|\pi_{\mathrm{ref}}]$ throughout training (Figure 6). Results show VinePPO consistently achieves higher accuracy at same KL divergence, indicating more efficient use of the “KL budget.” This efficiency stems from VinePPO’s more precise credit assignment. As shown in Figure 1, many advantages are zero, and VinePPO excludes these steps from the loss. By avoiding unnecessary updates on non-contributing tokens, VinePPO reduces non-essential parameter adjustments that would inflate KL. See Section D.1 for full results.
6.3 Temperature Tolerance
Sampling temperature is a critical hyperparameter controlling the randomness of sampled trajectories. At higher temperatures models generates more diverse trajectories, accelerating early training through increased exploration. However, this diversity challenges PPO’s value network, requiring generalization over a wider range of states. We compared VinePPO and PPO using temperatures $T\in\{0.6,0.8,1.0\}$ over the initial third of training steps. Figure 8 shows VinePPO consistently benefits from higher temperatures, achieving faster convergence. Conversely, PPO fails to benefit from increased exploration and even diverges at $T=1.0$, where trajectories are most diverse.
Figure 9: Distribution of predicted values for each state vs. ground truth (computed using 256 MC samples) during training for DeepSeekMath 7B on MATH dataset, highlighting the nature of errors. VinePPO achieves much lower Mean Absolute Error (MAE).
@@ 6.4 Computational Efficiency
VinePPO and PPO require different resources: PPO uses a separate value network, requiring two times more GPU memory (up to 112GB for a 7B LLM, considering both model and its optimizer); VinePPO, conversely, relies on MC samples. This skips value network’s memory requirements, but shifts the computational burden to increased LLM inferences, making VinePPO generally slower per iteration (up to 5x for RhoMath 1.1B and 2x for DeepSeekMath 7B). However, the effect of VinePPO’s accurate credit assignment is substantial. Although slower per iteration, VinePPO achieves PPO’s peak accuracy in fewer gradient steps and less wall-clock time. Figure 7 shows RhoMath 1.1B and DeepSeekMath 7B require about 3.0x and 1.51x less time and 9x and 2.8x fewer steps. This improvement occurs despite all hyperparameters being tuned for PPO. Therefore, switching to VinePPO offers a way to enhance performance within the same compute budget and serves as the only option when memory is constrained.
7 Value Prediction Analysis
In this section, we explore the underlying reasons for the performance gap between PPO and VinePPO by closely analyzing the value prediction of both methods. First, we establish a “ground truth” value at each reasoning step within trajectories by running many MC samples ($256$ in our case) and averaging the returns. This provides a low-variance reference value. We then compare the value predictions in both methods against this ground truth. We present the results for DeepSeekMath 7B on the MATH dataset (full analysis with other models and datasets in Section D.2).
@@ 7 Value Prediction Analysis
Figure 9 presents the distribution of value predictions at each reasoning step. The errors produced by VinePPO and PPO differ significantly. VinePPO’s estimates are unbiased, with variance peaking at 0.5 and dropping to zero at 0 and 1. PPO’s value network shows high bias, often misclassifying bad states (ground truth near 0) as good and vice versa. To further visualize accuracy, we classify a value prediction as “correct” if it falls within 0.05 of the ground truth. The accuracy of this formulation is shown in Figure 11.a. PPO’s value network starts with low accuracy, gradually improving to 65%. VinePPO, however, consistently achieves 70-90% accuracy throughout training.
@@
In value-based RL, ranking actions correctly is more crucial than absolute value accuracy. While PPO, as a policy gradient method, requires accurate value estimates to compute meaningful advantages, it is still a compelling question whether PPO’s value network, despite its bias, can maintain correct action ranking. To investigate, we sample five new next steps from the same initial state and evaluate if the method correctly identifies the resulting next state with the highest ground truth value. As shown in Figure 11.b, PPO’s value network performs near chance levels for much of the training, with slight improvements over time. In contrast, VinePPO consistently identifies the top action with high accuracy throughout training.
@@
To understand value computation mechanisms, we visualize the prediction error at each reasoning step within a trajectory. As shown in Figure 10, PPO’s estimation error increases as reasoning progresses. We hypothesize this occurs because early steps have lower diversity and resemble training data more, allowing the value network to rely on memorization. Later, as space of states become much larger, they become unfamiliar and the network struggles to generalize. VinePPO’s prediction error decreases with reasoning progression. We attribute this to the model becoming more deterministic in later steps as it conditions on bigger and longer context. This determinism enables more accurate estimates from the same number of MC samples.
Figure 10: Visualizing the Mean Absolute Error (MAE) of the value predictions at different point of the reasoning chain. Value Network in PPO fails to generalize as the reasoning chain progresses, while VinePPO’s value estimates become more accurate as the model become more deterministic.
Figure 11: (a) Value prediction accuracy formulated as a classification problem, where a prediction is considered correct if it falls within 0.05 of the ground truth. (b) Accuracy of identifying the top action in a set of five possible next states. VinePPO consistently outperforms the value network.

### v2

@@ 4 Accurate Credit Assignment with VinePPO
As outlined in Section 3, a step in the PPO gradient update aims to increase the probability of better-than-average actions while decreasing the probability of those that perform worse—a process quantified by the advantage $A(s_{t},a_{t})$. However, the true advantage is generally unknown and must be estimated, typically by substituting estimates from a value network into Section 3. As we will elaborate in Section 7, value networks are often inaccurate and result in biased value computation. Fortunately, the language environment as an MDP (Section 3) offers a useful property that allows for unbiased estimation of $V(s_{t})$. Since states are simply concatenated tokens, we can prompt the language model $\pi_{\theta}$ to generate continuations from any intermediate state. This flexibility allows us to explore alternative future paths from arbitrary points in a generation.
Specifically, computing the advantage requires access to $V(s_{t})=\mathbb{E}\left[R(\tau)\mid s_{0}=s_{t}\right].$ VinePPO obtain an MC estimation of this expectation by randomly sampling continuations and averaging their returns. That is, for each state $s_{t}$ in a training trajectory $\tau$, we utilize the resetting property and re-feed the partial context corresponding to $s_{t}$ to the current policy to sample $K$ auxiliary rollouts $\eta_{1},\dots,\eta_{K}\sim\pi_{\theta}(\cdot\mid s_{t})$. The empirical mean of returns across these rollouts serves as the value estimate:
$\hat{V}_{\mathrm{MC}}(s_{t})=\frac{1}{K}\sum_{k=1}^{K}R(\eta_{k}).$ (5)
Critically, $\eta_{k}$’s are used exclusively for value estimation and do not contribute directly to policy gradient updates as we lack CA on them. Once the value $\hat{V}_{\mathrm{MC}}(s_{t})$ is computed, we estimate the advantages of each action using Section 3:
$\hat{A}_{\mathrm{MC}}(s_{t},a_{t})=r(s_{t},a_{t})+\gamma\hat{V}_{\mathrm{MC}}(s_{t+1})-\hat{V}_{\mathrm{MC}}(s_{t}).$ (6)
For any $K\geq 1$, the policy gradient computed using the advantage estimator $\hat{A}_{\mathrm{MC}}$ is an unbiased estimate of the gradient of expected return ${\bm{g}}_{\mathrm{pg}}$. PPO framework then uses $\hat{A}_{\mathrm{MC}}$ to update the policy on trajectory $\tau$.
Variance and computational efficiency represent core trade-offs in every Monte Carlo estimation. Here, the sampling parameter $K$ control such tradeoff— increasing $K$ reduces estimator variance at the expense of increased sampling demands. In Section 6, we rigorously characterize these properties for VinePPO.
To enhance the efficiency of $\hat{A}_{\mathrm{MC}}$, we group states within a reasoning step and compute a single advantage, which is assigned to all tokens in that step (examples in Appendix B). This trades off granularity for efficiency, allowing finer resolution with more compute, or coarser estimates with limited resources. Furthermore, modern LLM inference engines (Kwon et al., 2023; Zheng et al., 2024) enable rapid on-the-fly generation [^4] Achieving up to 5K tokens/second on a single Nvidia A100 GPU for 7B LLMs in bfloat16., making our MC-based approach computationally practical at scale.
By restricting modifications only to the advantage computation stage of PPO, our approach also isolates the effects of improved credit assignment, revealing how unbiased advantage estimation fundamentally alters policy optimization dynamics compared to value-network baselines.
Figure 4: VinePPO outperforms standard PPO, GRPO, RLOO, and other RL-free baselines on Pass@1 performance on MATH and GSM8K datasets, while also exhibiting scalability across different model sizes.
Figure 5: Generalization slope improves with improved credit assignment. VinePPO has steepest generalization: making the highest generalization gains than baselines when fitting the same amount of training data. On the other end of CA spectrum, RestEM overfits its training data.
@@ 5 Experimental Setup
We conduct experiments on publicly available LLMs and datasets to ensure reproducibility. We use base versions of DeepSeekMath 7B (Shao et al., 2024) and RhoMath 1.1B (Lin et al., 2024) which are pretrained on mathematical and natural language corpora. We chose mathematical reasoning datasets MATH (Hendrycks et al., 2021), competition-level mathematical problems, and GSM8K (Cobbe et al., 2021), simpler grade-school level math word problems. Both datasets are well-established and present a range of difficulty levels. For each dataset, we finetune the base LLM on its respective training set to obtain the initial SFT policy ($\pi_{\mathrm{ref}}$). Throughout the paper, model names refer ones initialized from these SFT checkpoints. We employ full-parameter finetuning to leverage the models’ full capacity (Biderman et al., 2024).
Baselines
Our main baseline is the standard PPO framework (Ouyang et al., 2022; Huang et al., 2024), which VinePPO builds on and improves through better credit assignment. We also compare against PPO variants that forego the credit assignment: RLOO (Ahmadian et al., 2024) and GRPO (Shao et al., 2024). For RL-free alternatives, we include RestEM (Singh et al., 2024), a form of iterative rejection finetuning (Yuan et al., 2023; Anthony et al., 2017), and DPO+ (Pal et al., 2024), a working variant of DPO with strong performance on reasoning. Except VinePPO and standard PPO, all other baselines omit explicit credit assignment by design: i.e. they assign the same weight to all the tokens of a response. All methods use the same SFT checkpoint to ensure fair comparison. For each experiment, we choose the best checkpoint based on a held-out validation set. We compare all methods by accuracy (Pass@1) on test sets, measuring the correctness of final answers.
@@
We adopt a binary task reward $\mathcal{R}$ that evaluates final answer correctness against ground truth, following previous work (Pal et al., 2024; Singh et al., 2024). To ensure fair comparison, all methods consume the same number of episodes during training: for each question, we sample eight episodes and go over the dataset 8 times, yielding 64 episodes per question across all methods. For PPO, we first conduct an extensive hyperparameter search (such as KL penalty coefficient, batch size, minibatch size, GAE $\lambda$, number of epochs per iteration) and rigorously implement all established best practices and well-known techniques (Huang et al., 2024; Ivison et al., 2024) (Refer to Section C.2 for the full list). This ensures our evaluation reflects PPO’s state-of-the-art configuration and its full potential. VinePPO inherits PPO’s exact hyperparameters and only modifies the advantage estimation, keeping the rest unchanged. This design allows us to isolate the effect of refined credit assignment. For PPO variants (RLOO, GRPO), we closely follow their HuggingFace implementations. For these, we initialize with PPO’s hyperparameters but perform additional tuning to stabilize training while maintaining the same episode budget. For RL-free baselines (RestEM, DPO+), we strictly adhere to their original implementations (Singh et al., 2024; Pal et al., 2024) and match their sample consumption to other RL methods. For $\hat{V}_{\mathrm{MC}}$ in VinePPO, we conduct a full ablation study on $K$ in Section 6.1, with $K=9$ used as the default setting unless otherwise specified. To ensure a fair comparison of compute efficiency, we conduct controlled experiments in Section 6.2, where all methods are evaluated under identical hardware and parallelization protocols. Full implementation details, including hyperparameters and training procedures, are documented in Section C.6 to ensure reproducibility.
Figure 7: (a) Effect of the number of auxiliary rollouts $K$ for estimating $\hat{V}_{\mathrm{MC}}(s_{t})$ on RhoMath 1.1B and MATH (see Figure D.11 for GSM8K). Increasing $K$ consistently improves accuracy. (b) Wall-clock time for the same experiments. While increased sampling makes each iteration slower, the reduced variance leads to faster overall convergence.
Figure 8: (a) Visualizing the Mean Absolute Error (MAE) of the value predictions at different point of the reasoning chain. Value Network in PPO fails to generalize as the reasoning chain progresses, while VinePPO’s value estimates become more accurate as the model become more deterministic. (b) Accuracy of identifying the top action in a set of five possible next states. VinePPO consistently outperforms the value network.
@@ 6 Results
In this section, we evaluate the effect of better CA on task performance, efficiency, and generalization dynamics.
@@ 6.1 Task Performance
VinePPO consistently outperforms standard PPO throughout training (Figure C.4) and other baselines (Figure 4) achieving the highest test accuracy on both models and datasets. Notably, the performance gap widens in MATH which is more challenging than GSM8K. To confirm that PPO’s limitations are not due to undertrained value networks, we measured their explained variance, a standard metric for value function quality, which ranged between 0.7–0.9 across tasks (Figure D.5), indicating a well-trained critic. Because the PPO and VinePPO runs only differ in their value estimation, comparing these two isolates the effect of CA. As shown in Figure D.8, VinePPO reaches higher test accuracy given a limited KL budget. Additionally, VinePPO is more robust to higher sampling temperatures (Figure D.10).
@@ 6.2 Computational Efficiency
Training on a single trajectory in GRPO, RLOO, RestEM and DPO+, involves a forward and backward pass. PPO and VinePPO have extra computations of different types. PPO uses double GPU memory — the value network needs 112GB for a 7B LLM, considering both model and its optimizer. Additionally, PPO requires a forward pass for value prediction and a forward-backward pass for value network training. VinePPO replaces the value network with MC samples. Since generation is expensive, each step of VinePPO is slower (up to 5x for RhoMath 1.1B and 2x for DeepSeekMath 7B compared to PPO). VinePPO compensates for slower iterations by making each one more effective through better CA. Under the same hardware, it achieves higher test accuracy faster than baselines (Figure 6). Specifically, VinePPO matches PPO’s peak accuracy in fewer gradient steps and less wall-clock time. Figure 6 shows RhoMath 1.1B and DeepSeekMath 7B require about 3.0x and 1.51x less time and 9x and 2.8x fewer steps compared to PPO. This improvement occurs despite all hyperparameters being tuned for PPO. Therefore, switching to VinePPO could enhance the performance within the same compute budget.
6.3 Generalization Slope
High-quality and challenging reasoning tasks are scarce, making generalization a key challenge. Once a training instance is fitted, it provides no further signal for generalization. Thus, algorithms that maximize generalization efficiency are superior—achieving higher test accuracy for a given train accuracy. As shown in Figure 5, VinePPO demonstrates the strongest generalization gains compared to all other baselines. Notably, RestEM overfits near the end. This aligns with recent findings that RL generalizes while SFT primarily memorizes (Chu et al., 2025). Overall, allocating more compute to refining credit assignment, rather than brute-force data fitting, leads to stronger generalization.
6.4 Effect of $K$
We assess the impact of $K$, the number of MC samples, by running an ablation on RhoMath 1.1B, varying $K$ from 1 to 3 and 9. As shown in Figure 7, VinePPO improves with higher $K$ since more MC samples reduce the variance of $\hat{A}_{\mathrm{MC}}$. While high variance of MC estimation could theoretically hinder training, our results show that even small $K$ values work well in this setting. Interestingly, increasing $K$ also improves compute efficiency. Although each iteration takes longer, it becomes more effective. This suggests that increasing $K$ provides a practical way to leverage additional computational resources for better performance.
7 Why and How Value Networks Fail
In this section, we analyze the performance gap between PPO and VinePPO by focusing on their value predictions—their only difference. First, We establish a “ground truth” value at each reasoning step within trajectories by running $256$ MC samples and averaging the returns. Next, We compare the value predictions against this ground truth [^5] The return of sampled trajectory starting from a step is a bernoulli random variable. Let $\theta$ denote the true success probability. Our estimator, $\bar{X}=\frac{1}{256}\sum_{i=1}^{256}X_{i}$, has variance $\frac{\theta(1-\theta)}{256}$. At $\theta=0.5$ (maximal variance), this becomes $\frac{0.25}{256}\approx 0.001$.. We present the results for DeepSeekMath 7B , our biggest model, on the MATH dataset (all results in Section D.5).
@@ 7 Why and How Value Networks Fail
Figure 3 presents the distribution of value predictions at each reasoning step. VinePPO’s estimates are unbiased, with variance peaking at 0.5 and dropping to zero at 0 and 1. PPO’s value network shows high bias and often misclassifies bad states (ground truth near 0) as good and vice versa. We define a prediction as “correct” if it is within 0.05 of the ground truth. As shown in Figure D.12 PPO’s value network starts with low accuracy, gradually improving to 65%. In contrast, VinePPO consistently achieves 70-90% accuracy throughout training.
@@
In value-based RL, accurately ranking actions is more important than accurate value estimates. While PPO, a policy-based method, depends heavily on accurate value estimates, it raises an interesting question: Can PPO’s value network still rank actions correctly? We tested this by sampling five possible next steps from a shared initial state and measuring whether the method predicted the next step with the highest ground truth value by assigning it the highest predicted value. As shown in Figure 8.b, PPO’s value network performs near chance levels for most of the training, improving only slightly over time. In contrast, VinePPO consistently identifies the top action with high accuracy throughout.
@@
To understand value estimation dynamics, we plot value estimation error against reasoning step position (normalized; 3rd of 10 steps = 0.3). As shown in Figure 8.a, PPO performs worse as reasoning progresses. We hypothesize this is because early steps resemble training data, allowing the value network to rely on memorization. Later steps are more diverse and value network struggles to generalize. VinePPO’s prediction error decreases with reasoning progression. We attribute this to greater determinism in later steps, as the model conditions on a longer context. This stability improves value estimation from the same number of MC samples.